# Introducing NewCluster: the first half of the history of a high-resolution cluster simulation

### v1

@@ 3.1 Overview of the Simulation
We also present the mock images of randomly selected galaxies at three different redshifts (1.5, 1.15, and 0.8) in Fig. 2. We used the Sloan Digital Sky Survey (SDSS) $r$-/$g$-/$u$-band fluxes to create the RGB composite images. We can clearly notice the diversity of galaxy morphologies in NewCluster  especially with attenuation effects from the on-the-fly calculated dust.
@@ 3.2 Galaxy Scaling Relations
In Fig. 3, we present some key scaling relations of galaxies from the NewCluster simulation at four different redshifts, $z=0.8,1.5,2.5$, and $4.0$. Each panel shows different galaxy properties on the y-axis as a function of their stellar mass on the x-axis. The black lines indicate empirical relations from observations, while the background pixels show the distribution of NewCluster galaxies at $z=0.8$. The circles indicate median values of NewCluster galaxies at each snapshot, the standard error of the mean is shown as vertical lines only for $z=0.8$ snapshot. Panel (a) shows the SFR versus stellar mass relation in comparison with observed main sequence from Whitaker et al. (2014). The distribution of galaxies at $z\sim 0.8$ from Laigle et al. (2016) is also shown with $1\sigma$ and $2\sigma$ distributions as solid and dotted contours, respectively. SFR is measured based on the stars that are recognised as members of each galaxy. An overall decrease in the SFR sequence of NewCluster galaxies is observed with decreasing redshift. The sequence at $z=0.8$ shows SFR lower than the empirical relation. This could be attributed to the fact that the zoom-in region lies in a highly over-dense environment, where galaxies form earlier, and environmental quenching becomes effective. As a result, the galaxy population is significantly older than that of the average field.
@@
Panel (b) shows the fraction of cold gas in galaxies as a function of stellar mass. The cold gas mass is measured by summing the gas mass of cells with temperatures below $10^{4}\,\rm{K}$ within a radius of $2\,r_{50}$ from the galactic centre. Each median curve for the simulated galaxies has been measured for galaxies containing cold gas. The observational relations are from Catinella et al. (2018) and Zhang et al. (2021), where the former measured both atomic and molecular hydrogen at $z=0$, while the latter measured only atomic hydrogen gas at $z=0.8$. The gas mass of NewCluster galaxies does not change significantly with redshift, which is consistent with the result of Dubois et al. (2021). However, we find that the fraction of galaxies without any cold gas changes significantly with redshift, as shown in the embedded figure in the panel. This indicates that gas removal occurs in certain galaxies through the violent depletion of cold gas, likely driven by environmental effects.
@@
Panel (c) shows galaxy sizes, represented by the half mass radius $r_{50}$. To obtain a uniform distribution of lines of sight, we used 100 directions generated from Fibonacci lattices to measure apparent radii and used the median value of the distribution. Empirical relations are shown from Mowla et al. (2019), which includes all galaxy types (solid line), and from Nedkova et al. (2021), which separates galaxies into quiescent (dashed line) and star-forming (dotted line) types. The sizes of NewCluster galaxies show a moderate increase over time for lower-mass galaxies. The decrease in size-mass relation at high-redshift could be attributed to the gas-rich compaction event that takes place on galaxies above $\sim 10^{10}\,M_{\sun}$ (Dekel & Burkert, 2014; Zolotov et al., 2015; Lapiner et al., 2021). These high-mass galaxies later show strong redshift evolution toward larger radii, likely driven by mergers (van der Wel et al., 2014). At $z=0.8$, the empirical relation of star-forming galaxies closely overlaps with the lower-mass sequence of the simulated galaxies in the range of $10^{8}$$10^{10}\,\,M_{\sun}$. The consistency seems to degrade at higher masses, while a plateau appears in the high-mass end of $10^{10}-10^{11}\,M_{\sun}$. This can also be interpreted as the population transition from extended star-forming galaxies to compact, quiescent galaxies with increasing mass, where quenching is induced by the combination of mass and environmental effects within the cluster. Simulated galaxies more massive than $10^{11}\,M_{\sun}$ show smaller sizes compared to the empirical relation. This discrepancy is mainly due to their overly compact central stellar concentrations, which significantly reduce their half-mass radii, even though the most massive galaxies in the simulation clearly extend well beyond $\sim 10\,\rm{kpc}$ in radius.
@@
Panel (d) presents the stellar mass versus stellar metallicity relation of NewCluster galaxies, measured in terms of [O/H], in comparison with empirical relations from Savaglio et al. (2005) and Wuyts et al. (2014). The NewCluster galaxies show no strong indication of evolution in the mass–metallicity relation across redshift. At $z=0.8$, simulated galaxies show a tight relation that is consistent with observations, representing the continuous enrichment and recycling of gas within galaxies along with star formation. A minor population of galaxies appears in the upper-left region of the main sequence. Most of them turn out to be satellite galaxies that have undergone severe stripping, resulting in stellar mass loss that causes them to deviate from the main relation.
@@
Panel (e) shows the stellar rotational velocity $V$ divided by the velocity dispersion $\sigma$ within galaxies. NewCluster galaxies exhibit a pronounced evolution in the relation across redshift. At high redshifts ($z=2.5$ and $4.0$), galaxies show stronger rotation at higher masses, consistent with a disc settling scenario that preferentially occurs in massive galaxies (Park et al., 2019). At low redshifts ($z=0.8$ and $1.5$), the most massive galaxies evolve into less rotationally supported systems, consistent with the formation of massive slow rotators through continuous mergers (Gerhard, 1981; Barnes, 1988; Choi & Yi, 2017). This behaviour is consistent with Dubois et al. (2016), who found that the $(V/\sigma)_{*}$ trend peaks around a stellar mass of $10^{11}\,\,M_{\sun}$. The median relation derived from the observed sample of Falcón-Barroso et al. (2019) at $z\sim 0$ is shown as a black dashed line, which follows a similar trend to the extrapolated redshift evolution of the simulated galaxies toward lower redshift.
@@
Panel (f) shows the mass of the most massive black holes within each galaxy, compared with the empirical relations of Suh et al. (2020) and Reines & Volonteri (2015). The former is derived from observations of broad-line AGNs at $z<2.5$, shown as the solid line, while the latter is derived from classical bulges and ellipticals, shown as the dashed line. Regardless of the redshift, NewCluster galaxies lie between the elliptical and AGN relations and exhibit a large scatter around the empirical trends, reflecting the morphological diversity of the sample as seen in Fig. 2. NewCluster galaxies show a moderate increase in MBH mass at fixed stellar mass with decreasing redshift, suggesting a possible redshift evolution in the black hole mass–host galaxy mass relation. However, this trend may also be influenced by biases arising from the general morphological evolution of galaxies, as early-type galaxies become increasingly common at lower redshifts.
@@
Figure 4 shows the stellar-to-halo mass ratio as a function of halo mass for NewCluster galaxies. The halo mass is measured at the peak mass of the main progenitor. Circles indicate the median of the distribution in each bin, colour-coded with different redshifts. The $1\sigma$ scatter at $z=0.8$ is shown as shades. Stars indicate individual galaxies with halo mass above $10^{13}\,M_{\sun}$. Empirical relations are shown for Behroozi et al. (2013a) as the solid line and Moster et al. (2010) as the dashed line. Although NewCluster galaxies lie in a high-density environment that may exhibit a stellar-to-halo mass ratio different from the universal relation, the excess of stellar mass in the intermediate-mass halos ($M_{\rm halo}\sim 10^{11}\,M_{\sun}$) may indicate the over-production of stars for dwarf galaxies. This suggests that the current stellar feedback processes are not effective at suppressing star formation in low-mass galaxies at high redshifts. However, their effects at $z<1$ appear to lead to an overly quenched population of galaxies, as shown in Panel (a) of Fig. 3.
@@
Figure 5: Photometric properties of NewCluster galaxies at $z=0.8$. The x-axis shows the $r$-band absolute magnitude. The y-axis of the top (bottom) panel represents $g-r$ colour ($g$-band effective surface brightness). We also present the black arrow to show the effect of dust attenuation assuming $\rm E(B-V)=0.1$ with Calzetti attenuation curve (Calzetti et al., 2000). The colour of data points indicates the mass-weighted age of galaxies. The grey contours in both panels represent local observations from SDSS (Nair & Abraham, 2010).
@@
We present the photometric properties of NewCluster galaxies at $z=0.8$ in Figure 5. The sample galaxies are the same as those in Fig. 3, but excluding misidentified clump-like substructures. We measure the $g$- and $r$-band absolute magnitudes using stellar age and metallicity, based on the simple stellar population models from Bruzual & Charlot (2007). [^5] We simply measure the photometric magnitudes from the averaged age and metallicity of stars following the table, not considering dust absorption. The top panel shows the $g-r$ colour versus $r$-band absolute magnitude, while the bottom panel presents the $g$-band surface brightness within the effective radius. Different colours indicate the $r$-band-weighted age of galaxies.
@@
The observed distributions of local galaxies are also displayed (Nair & Abraham, 2010). Although the overall and median (green error bar) locations are not identical to observations, the trends are consistent with naive expectations. For instance, our colour-magnitude distribution shows the separation between old and young populations, while the red sequence is not clearly separated at this redshift.
@@
In the bottom panel, central surface brightness also linearly correlates well with absolute magnitude, as reported in observations of the Local Group (McConnachie, 2012). We note that a separate cluster of objects is observed in the upper left of the main sequence. Their old age, low luminosity, and high surface brightness indicate that these objects are compact and quenched. After tracing their evolution through the merger tree, we find that they are not transient structures, as they show long, continuous histories without breaks. Since their progenitors are usually found as substructures within larger galaxies in terms of spatial distribution, they are likely ultra-compact dwarf galaxies (UCD) or orphan stellar systems that were ejected from larger galaxies (e.g., Jang et al., 2024). We leave the detailed investigation of these galaxies to future work.
@@
Figure 6 shows the distribution of the stellar alpha enhancement fraction ([Mg/Fe]) as a function of metallicity ([Fe/H]). The grey background indicates the entire population of stars in the simulation, while the two contours show the members of galaxies in two different stellar mass ranges. As reported by previous studies (e.g., Bensby et al., 2014), the alpha‐element enhancement fraction decreases with increasing metallicity, reflecting the differential enrichment from SNII and SNIa with their respective delay‐time distributions. Comparing different stellar mass ranges, the stellar populations in more massive galaxies show overall higher metallicities and, at a given metallicity, higher alpha-element enhancement fractions. This indicates that the more massive galaxies in our simulation had more recycling of gas but with a shorter star formation history.
@@ 3.3 Low-surface Brightness Structures
Low-surface brightness (LSB) structures, such as tidal streams, UDGs, and ICL, are emerging as key tracers of dynamical evolution through the assembly history (van Dokkum et al., 2015; Martin et al., 2019; Contini, 2021; Contini et al., 2023). They are observed on a wide range of scales, from dwarf galaxies to massive galaxy clusters, and are generally defined as regions fainter than a surface brightness threshold of $22-24\,\rm{mag\,arcsec^{-2}}$ (Bothun et al., 1997; Bakos & Trujillo, 2012; Tanoglidis et al., 2021). Thanks to the fine stellar mass resolution of our simulation, we are able to resolve these faint structures at $z\sim 0.8$. LSB features of galaxies are receiving increasing attention due to their potential to shed light on the past assembly history of galaxies. UDGs are also interesting targets for studying extreme sites of star formation and their distinct characteristics compared to typical dwarf galaxies (van Dokkum et al., 2015; Pandya et al., 2018; Kado-Fong et al., 2022). Moreover, ICL has the potential to serve as an observable tracer of a cluster’s dynamical evolution and even the distribution of DM (Montes & Trujillo, 2019; Contini, 2021; Contini et al., 2023).
@@
Despite being a broadly appealing target for study, LSB structures are inherently difficult to investigate due to their low detectability (Contini, 2021; Yi et al., 2022; Joo & Jee, 2023; Kelvin et al., 2023; Kim et al., 2023a). This is especially true at high redshifts, where galaxies may exhibit more dramatic tidal features because they are dynamically younger (Hopkins et al., 2010; Deger et al., 2018; Ferreira et al., 2020), but most studies are based on observations at $z<1$ due to detection limits reduced by cosmological dimming. The emergence of cutting-edge instruments (e.g., JWST, LSST) presents an opportunity to extend our studies to higher redshifts and is expected to bring numerous LSB detections. It is necessary to prepare corresponding high-resolution simulations to resolve faint structures in a cosmological context in order to capture a wide range of their formation scenarios.
@@
The minimum stellar particle mass of NewCluster is $\sim 2\times 10^{4}\,M_{\odot}$, which corresponds to a theoretical $r$-band surface brightness limit of $33.5\,\rm{mag\,arcsec^{-2}}$ at $z=0.8$. This allows us to resolve faint structures on various scales. In the top panels of Fig. 7, we display stellar density maps of the ICL, a tidally distorted galaxy, and an LSB galaxy from left to right. Many faint structures are visible, not only in the ICL, but also in the trails of satellite galaxies. As more dynamical interactions are expected among galaxies in cluster environments (Dressler, 1980; Jaffé et al., 2011; Yi et al., 2013), our simulation is well suited for studying the evolution of interacting galaxies down to the dwarf galaxy regime.
@@
To enable direct comparison between our model and high-redshift observations, we also generate mock SDSS- and JWST-like images for these targets, shown in the middle and bottom panels of Fig. 7.666Since the wavelength coverages of SDSS and JWST are different, we arbitrarily select filters of $r$, $i$, and $z$ bands of SDSS, which correspond to nearly the F070W and F090W filters of JWST. Point spread functions (PSFs) at these JWST wavelengths are generated using the STPSF package (formerly WebbPSF; Perrin et al., 2012). We note that the mock SDSS images do not include the atmospheric seeing effect and therefore represent an idealised case compared to actual SDSS observations. It is challenging to capture the LSB features at $z\sim 0.8$ with SDSS-like instruments (middle row). In particular, LSB galaxy is not visible in the SDSS mock image due to limited resolution and sensitivity, even though we select a relatively massive ($M_{*}=3.5\times 10^{9}\,M_{\sun}$) galaxy. In contrast, applying the resolution and PSF of JWST (bottom row) enables the detection of faint structures seen in the stellar density map (top row).

### v2

@@ 3.1 Overview of the simulation
We also present the mock images of randomly selected galaxies at three different redshifts (1.5, 1.15, and 0.8) in Fig. 2. We used the Sloan Digital Sky Survey (SDSS) rr-/gg-/$u$-band fluxes to create the RGB composite images. We can clearly notice the diversity of galaxy morphologies in NewCluster  especially with attenuation effects from the on-the-fly calculated dust.
@@ 3.2 Galaxy scaling relations
In Fig. 3, we present some key scaling relations of galaxies from the NewCluster simulation at four different redshifts, $z=0.8,1.5,2.5$, and $4.0$. Each panel shows different galaxy properties on the $y$-axis as a function of their stellar mass on the $x$-axis. The black lines indicate empirical relations from observations, while the background pixels show the distribution of NewCluster galaxies at $z=0.8$. The circles indicate median values of NewCluster galaxies at each snapshot, the standard error of the mean is shown as vertical lines only for $z=0.8$ snapshot. Panel (a) shows the SFR versus stellar mass relation in comparison with observed main sequence from 2014ApJ...795..104W. The distribution of galaxies at $z\sim 0.8$ from 2016ApJS..224...24L is also shown with $1\sigma$ and $2\sigma$ distributions as solid and dotted contours, respectively. SFR is measured based on the stars that are recognised as members of each galaxy. An overall decrease in the SFR sequence of NewCluster galaxies is observed with decreasing redshift. The sequence at $z=0.8$ shows SFR lower than the empirical relation. This could be attributed to the fact that the zoom-in region lies in a highly over-dense environment, where galaxies form earlier, and environmental quenching becomes effective. As a result, the galaxy population is significantly older than that of the average field.
@@
Panel (b) shows the fraction of cold gas in galaxies as a function of stellar mass. The cold gas mass is measured by summing the gas mass of cells with temperatures below $10^{4}\,\rm{K}$ within a radius of $2\,R_{50,3D}$ from the galactic centre. Each median curve for the simulated galaxies has been measured for galaxies containing cold gas. The observational relations are from 2018MNRAS.476..875C and 2021A&A...648A..25Z, where the former measured both atomic and molecular hydrogen at $z=0$, while the latter measured only atomic hydrogen gas at $z=0.8$. The gas mass of NewCluster galaxies does not change significantly with redshift, which is consistent with the result of 2021A&A...651A.109D. However, we find that the fraction of galaxies without any cold gas changes significantly with redshift, as shown in the embedded figure in the panel. This indicates that gas removal occurs in certain galaxies through the violent depletion of cold gas, likely driven by environmental effects.
@@
Panel (c) shows galaxy sizes, represented by the half mass radius in 2-dimension, $r_{\rm 50,2D}$. To obtain a uniform distribution of lines of sight, we used 100 directions generated from Fibonacci lattices to measure apparent radii and used the median value of the distribution. Empirical relations are shown from 2019ApJ...872L..13M, which includes all galaxy types (solid line), and from 2021MNRAS.506..928N, which separates galaxies into quiescent (dashed line) and star-forming (dotted line) types. The sizes of NewCluster galaxies show a moderate increase over time for lower-mass galaxies. The decrease in size-mass relation at high-redshift could be attributed to the gas-rich compaction event that takes place on galaxies above $\sim 10^{10}\,M_{\sun}$ (2014MNRAS.438.1870D; 2015MNRAS.450.2327Z; 2021MNRAS.505..172L). These high-mass galaxies later show strong redshift evolution toward larger radii, likely driven by mergers (2014ApJ...788...28V). At $z=0.8$, the empirical relation of star-forming galaxies closely overlaps with the lower-mass sequence of the simulated galaxies in the range of $^{8}1010$$^{10}\,\,M_{\sun}$. The consistency seems to degrade at higher masses, while a plateau appears in the high-mass end of $10^{10}-10^{11}\,M_{\sun}$. This can also be interpreted as the population transition from extended star-forming galaxies to compact, quiescent galaxies with increasing mass, where quenching is induced by the combination of mass and environmental effects within the cluster. Simulated galaxies more massive than $10^{11}\,M_{\sun}$ show smaller sizes compared to the empirical relation. This discrepancy is mainly due to their overly compact central stellar concentrations, which significantly reduce their half-mass radii, even though the most massive galaxies in the simulation clearly extend well beyond $\sim 10\,\rm{kpc}$ in radius.
@@
Panel (d) presents the stellar mass versus stellar metallicity relation of NewCluster galaxies, measured in terms of [O/H], in comparison with empirical relations from 2005ApJ...635..260S and 2014ApJ...789L..40W. The NewCluster galaxies show no strong indication of evolution in the mass–metallicity relation across redshift. At $z=0.8$, simulated galaxies show a tight relation that is consistent with observations, representing the continuous enrichment and recycling of gas within galaxies along with star formation. A minor population of galaxies appears in the upper-left region of the main sequence. Most of them turn out to be satellite galaxies that have undergone severe stripping, resulting in stellar mass loss that causes them to deviate from the main relation.
@@
Panel (e) shows the stellar rotational velocity $V$ divided by the velocity dispersion $\sigma$ within galaxies. NewCluster galaxies exhibit a pronounced evolution in the relation across redshift. At high redshifts ($z=2.5$ and $4.0$), galaxies show stronger rotation at higher masses, consistent with a disc settling scenario that preferentially occurs in massive galaxies (2019ApJ...883...25P). At low redshifts ($z=0.8$ and $1.5$), the most massive galaxies evolve into less rotationally supported systems, consistent with the formation of massive slow rotators through continuous mergers (1981MNRAS.197..179G; 1988ApJ...331..699B; CY17). This behaviour is consistent with Dubois16, who found that the $(V/\sigma)_{*}$ trend peaks around a stellar mass of $10^{11}\,\,M_{\sun}$. The median relation derived from the observed sample of 2019A&A...632A..59F at $z\sim 0$ is shown as a black dashed line, which follows a similar trend to the extrapolated redshift evolution of the simulated galaxies toward lower redshift.
@@
Panel (f) shows the mass of the most massive black holes within each galaxy, compared with the empirical relations of 2020ApJ...889...32S and 2015ApJ...813...82R. The former is derived from observations of broad-line AGNs at $z<2.5$, shown as the solid line, while the latter is derived from classical bulges and ellipticals, shown as the dashed line. Regardless of the redshift, NewCluster galaxies lie between the elliptical and AGN relations and exhibit a large scatter around the empirical trends, reflecting the morphological diversity of the sample as seen in Fig. 2. NewCluster galaxies show a moderate increase in MBH mass at fixed stellar mass with decreasing redshift, suggesting a possible redshift evolution in the black hole mass–host galaxy mass relation. However, this trend may also be influenced by biases arising from the general morphological evolution of galaxies, as early-type galaxies become increasingly common at lower redshifts.
@@
Figure 4 shows the stellar-to-halo mass ratio as a function of halo mass for NewCluster galaxies. The halo mass is measured at the peak mass of the main progenitor. Circles indicate the median of the distribution in each bin, colour-coded with different redshifts. The $1\sigma$ scatter at $z=0.8$ is shown as shades. Stars indicate individual galaxies with halo mass above $10^{13}\,M_{\sun}$. Empirical relations are shown for 2013ApJ...770...57B as the solid line and 2010ApJ...710..903M as the dashed line. Although NewCluster galaxies lie in a high-density environment that may exhibit a stellar-to-halo mass ratio different from the universal relation, the excess of stellar mass in the intermediate-mass halos ($M_{\rm halo}\sim 10^{11}\,M_{\sun}$) may indicate the over-production of stars for dwarf galaxies. This suggests that the current stellar feedback processes are not effective at suppressing star formation in low-mass galaxies at high redshifts. However, their effects at $z<1$ appear to lead to an overly quenched population of galaxies, as shown in Panel (a) of Fig. 3.
@@
Figure 5: Photometric properties of NewCluster galaxies at $z=0.8$. The $x$-axis shows the $r$-band absolute magnitude. The $y$-axis of the top (bottom) panel represents grg-r colour ($g$-band effective surface brightness). We also present the black arrow to show the effect of dust attenuation assuming $\rm E(B-V)=0.1$ with Calzetti attenuation curve (Calzetti2000). The colour of data points indicates the mass-weighted age of galaxies. The grey contours in both panels represent local observations from SDSS (Nair2010).
@@
We present the photometric properties of NewCluster galaxies at $z=0.8$ in Figure 5. The sample galaxies are the same as those in Fig. 3, but excluding misidentified clump-like substructures. We measure the $g$- and $r$-band absolute magnitudes using stellar age and metallicity, based on the simple stellar population models from CB07.555We simply measure the photometric magnitudes from the averaged age and metallicity of stars following the table, not considering dust absorption. The top panel shows the grg-r colour versus $r$-band absolute magnitude, while the bottom panel presents the $g$-band surface brightness within the effective radius. Different colours indicate the $r$-band-weighted age of galaxies.
@@
The observed distributions of local galaxies are also displayed (Nair2010). Although the overall and median (green error bar) locations are not identical to observations, the trends are consistent with naive expectations. For instance, our colour-magnitude distribution shows the separation between old and young populations, while the red sequence is not clearly separated at this redshift.
@@
In the bottom panel, central surface brightness also linearly correlates well with absolute magnitude, as reported in observations of the Local Group (McConnachie2012). We note that a separate cluster of objects is observed in the upper left of the main sequence. Their old age, low luminosity, and high surface brightness indicate that these objects are compact and quenched. After tracing their evolution through the merger tree, we find that they are not transient structures, as they show long, continuous histories without breaks. Since their progenitors are usually found as substructures within larger galaxies in terms of spatial distribution, they are likely ultra-compact dwarf galaxies (UCD) or orphan stellar systems that were ejected from larger galaxies (e.g., Jang2024). We leave the detailed investigation of these galaxies to future work.
@@
Figure 6 shows the distribution of the stellar alpha enhancement fraction ([Mg/Fe]) as a function of metallicity ([Fe/H]). The grey background indicates the entire population of stars in the simulation, while the two contours show the members of galaxies in two different stellar mass ranges. As reported by previous studies (e.g., 2014A&A...562A..71B), the alpha‐element enhancement fraction decreases with increasing metallicity, reflecting the differential enrichment from SNII and SNIa with their respective delay‐time distributions. Comparing different stellar mass ranges, the stellar populations in more massive galaxies show overall higher metallicities and, at a given metallicity, higher alpha-element enhancement fractions. This indicates that the more massive galaxies in our simulation had more recycling of gas but with a shorter star formation history.
@@ 3.3 Low-surface brightness structures
Low-surface brightness (LSB) structures, such as tidal streams, UDGs, and ICL, are emerging as key tracers of dynamical evolution through the assembly history (vanDokkum2015; Martin2019; Contini2021; Contini2023). They are observed on a wide range of scales, from dwarf galaxies to massive galaxy clusters, and are generally defined as regions fainter than a surface brightness threshold of $22-24\,\rm{mag\,arcsec^{-2}}$ (Bothun1997; Bakos2012; Tanoglidis2021). Thanks to the fine stellar mass resolution of our simulation, we are able to resolve these faint structures at $z\sim 0.8$. LSB features of galaxies are receiving increasing attention due to their potential to shed light on the past assembly history of galaxies. UDGs are also interesting targets for studying extreme sites of star formation and their distinct characteristics compared to typical dwarf galaxies (vanDokkum2015; Pandya2018; KadoFong2022). Moreover, ICL has the potential to serve as an observable tracer of a cluster’s dynamical evolution and even the distribution of DM (Montes2019; Contini2021; Contini2023).
@@
Despite being a broadly appealing target for study, LSB structures are inherently difficult to investigate due to their low detectability (Contini2021; Yi2022; Joo2023; Kelvin2023; KimJ23). This is especially true at high redshifts, where galaxies may exhibit more dramatic tidal features because they are dynamically younger (Hopkins2010; Deger2018; Ferreira2020), but most studies are based on observations at $z<1$ due to detection limits reduced by cosmological dimming. The emergence of cutting-edge instruments (e.g., JWST, LSST) presents an opportunity to extend our studies to higher redshifts and is expected to bring numerous LSB detections. It is necessary to prepare corresponding high-resolution simulations to resolve faint structures in a cosmological context in order to capture a wide range of their formation scenarios.
@@
The minimum stellar particle mass of NewCluster is $\sim 2\times 10^{4}\,M_{\odot}$, which corresponds to a theoretical $r$-band surface brightness limit of $33.5\,\rm{mag\,arcsec^{-2}}$ at $z=0.8$. This allows us to resolve faint structures on various scales. In the top panels of Fig. 7, we display stellar density maps of the ICL, a tidally distorted galaxy, and an LSB galaxy from left to right. Many faint structures are visible, not only in the ICL, but also in the trails of satellite galaxies. As more dynamical interactions are expected among galaxies in cluster environments (Dressler1980; Jaffe2011; Yi2013), our simulation is well suited for studying the evolution of interacting galaxies down to the dwarf galaxy regime.
@@
To enable direct comparison between our model and high-redshift observations, we also generate mock SDSS- and JWST-like images for these targets, shown in the middle and bottom panels of Fig. 7.666Since the wavelength coverages of SDSS and JWST are different, we arbitrarily select filters of $r$, $i$, and $z$ bands of SDSS, which correspond to nearly the F070W and F090W filters of JWST. Point spread functions (PSFs) at these JWST wavelengths are generated using the STPSF package (formerly WebbPSF; Perrin2012). We note that the mock SDSS images do not include the atmospheric seeing effect and therefore represent an idealised case compared to actual SDSS observations. It is challenging to capture the LSB features at $z\sim 0.8$ with SDSS-like instruments (middle row). In particular, LSB galaxy is not visible in the SDSS mock image due to limited resolution and sensitivity, even though we select a relatively massive ($M_{*}=3.5\times 10^{9}\,M_{\sun}$) galaxy. In contrast, applying the resolution and PSF of JWST (bottom row) enables the detection of faint structures seen in the stellar density map (top row).